# The Male CEO and the Female Assistant: Probing Gender Biases in Text-To-Image Models Through Paired Stereotype Test

### v1

@@ 4.5.2 Biases in Gendered Occupation
Figure 3: Micro STS under the PST setting by occupations, sorted in ascending order.
Micro STS Figure 3 visualizes the results of Micro STS for biases in gendered occupation, stratified by different professions. While noticeable levels of gender biases are found in all occupations, images generated for professions such as sheriff, mechanician, and secretary tend to carry the highest level of biases. Full quantitative results of the Micro STS for gendered occupation biases are provided in Appendix F.2, Table 14.
Figure 4: Visualization of how gender biases under PST align with real-world gender segregation for different occupations. Each pair of bars in the figure demonstrate the percentage of feminine figures in generated images, and the ratio of females in real-world labor statistics for a specific profession.
Next, we explore how results of PST demonstrate gender bias patterns that align with real-world biases. To more intuitively demonstrate this alignment of gender biases in DALLE-3 and in human society, we visualize the percentage of female individuals in the generated images and in the real world for each profession. Figure 4 shows how the level of gender imbalance in the generated images generally aligns with the gender segregation level in labor statistics. For socially male-stereotyped occupations such as sheriff and mover, DALLE-3 also generates a significantly smaller percentage of feminine individuals than for female-stereotyped professions like secretary and sewist.
Single-Person Setting vs. PST We conduct further analysis on how PST can reveal underlying patterns of gender biases, compared to the single-person setting. Figure 5 visualizes bias score differences between single-person and PST settings for all occupations. Interestingly, the $5$ occupations with the lowest bias level differences are all female-stereotyped, including “writer”, “clerk”, “auditor”, “baker”, and “designer”.
Figure 5: Visualization of differences between the level of occupational gender biases in paired setting vs. single-person setting, by occupations.
Figure 6: Detailed illustration of how gender biases in organizational power manifest under the proposed PST setting, when images generated in the single-person setting seem unbiased. The green “-” sign indicates anti-stereotype, whereas the red “+” sign highlights gender-stereotyped images.
Occupation Category Micro STS (Single)normal-$\downarrow$ Micro STS (Paired)normal-$\downarrow$ Micro STS (Gap)
Overall
$10.00$
47.38
$37.38$
Female-stereotyped
50.00
$45.00$
$-5.00$
Male-stereotyped
$-30.00$
49.74
$79.74$
Table 3: STS Differences for bias in gendered occupation, averaged over female-stereotypical and male-stereotypical occupations.
@@ 4.6 Biases in Organizational Power
Micro STS We explore how gender biases in stereotypical organizational power dynamics may manifest under the PST setting. Figure 7 visualizes the organizational power bias scores of “powerful” and “powerless” positions. It can be observed that while a noticeable amount of gender biases exist in organizational power dynamics under the PST setting, powerful positions possess a higher average level of biases. This means that DALLE-3 has a higher tendency to depict individuals of male-stereotyped powerful positions as being masculine. The observation aligns with previous conclusions on biases in gendered occupation, for which DALLE-3 also tends to generate masculine individuals for male-stereotypical occupations.
Figure 7: Visualization of organizational power bias levels under PST. Each marker indicates the level of bias for one occupation.
Power Level Micro STS (Single)normal-$\downarrow$ Micro STS (Paired)normal-$\downarrow$ Micro STS (Gap)
Overall
$4.62$
18.98
$14.36$
Powerless
37.04
$17.60$
$-19.44$
Powerful
$-27.78$
20.38
$48.16$
Table 4: Differences between Micro STS for biases in organizational power of the single-person setting vs. the proposed paired setting.
Single-Person Setting vs. PST Quantitative results in Table 16 further prove the above-mentioned observation. Additionally, while the PST framework unveils a higher overall level of organizational power bias, the single-person setting is more biased towards generating feminine individuals for powerless positions. We additionally visualize bias score differences between single-person and PST settings in Figure 8. The $2$ different markers represent male-stereotypical and female-stereotypical occupations. We found that powerful positions in female-stereotypical occupations are notably more likely to be depicted as masculine figures in the PST setting than the single-person setting, resulting in a majority of positive bias score differences.
Figure 8: Visualization of differences between the level of organizational power gender biases under the PST setting vs. in the single-person setting.
5 Mitigating Gender Biases in DALLE-3
@@
Given that the proposed PST setting reveals additional aspects of gender biases in T2I models, a natural following question would be “How to mitigate these biases?”. Since our experiments are all on DALLE-3, a completely black-box model, only prompt-based methods can be applied to reduce biases. Inspired by  Bansal et al. (2022)’s work, we conducted an exploration of how the use of textual fairness interventions can influence the level of biases in models. We took one of the fairness-related intervention prompts in  Bansal et al. (2022)’s study and adapted the prompt for the paired generation setting in PST. The fairness-intervened prompt is provided in Table 5.
Fairness Intervened Prompt
@@
Generate a picture of one {occ1} on the left and one {occ2} on the right, if all individuals can be a {occ1} or a {occ2} irrespective of their gender.
Table 5: Prompt for fairness-intervened T2I generation.
@@
Experiment results in Table 6 demonstrate the outstanding performance of the fairness intervention method for bias mitigation. Note that for biases in gendered occupation, we use fairness intervention to re-generated images with the 5 most biased occupations for males and females, respectively. The overall level of biases, as measured by the STS score, drops significantly after applying the interventions for both aspects of gender bias. However, the fairness-intervened image generation method seems to overdress biases in organizational power, resulting in a slight bias in the opposite direction. Controllability and interpretability of bias mitigation methods still remain an important research question in future studies.
Bias Aspect Mitigation Overall STSnormal-$\downarrow$
Gendered Occupation (Top5)
None
$92.00$
Intervention
26.00
Organizational Power
None
$18.98$
Intervention
-11.12
@@
6 Conclusion
This study is amongst the first to propose the Paired Stereotype Test (PST) setting to uncover more complicated patterns of gender biases in Text-To-Image (T2I) models. PST prompts T2I systems to generate two individuals at the same time, while they are described with two different social identities that are stereotypically associated with the opposite gender. By assessing the level of gender stereotype conformation in model generations under this paired setting, the PST framework is able to probe for more underlying gender biases in models than the previously-used single-person evaluation setting. Through experimenting on the DALLE-3 model, we use PST to reveal gender biases in the T2I system from two aspects: biases in gendered occupation and biases in organizational power. For biases in gendered occupation, we found that DALLE-3 is heavily biased toward generating masculine figures for individuals with male-stereotypical occupations, and generating feminine figures for those with female-stereotypical professions. We also provide interesting insights on how the level of biases in gendered occupations in DALLE-3’s generations aligns with real-world gender segregation statistics for the corresponding professions. For biases in organizational power, we discovered that DALLE-3 tend to portray individuals with powerful organizational positions like “manager” to be masculine, whereas depicting those with powerless positions as feminine. Furthermore, we show that while DALLE-3’s generations under the single-person setting are seemingly unbiased, the proposed PST setting effectively reveals the underlying patterns of gender biases. Our work highlights the severity and complexity of the issue of gender biases in T2I systems, as well as provides novel methodologies for bias evaluation that directly contribute to future fairness studies along this direction.
@@ 6 Conclusion
We identify the major limitations of our study. Firstly, since there lack previous literature on gender biases towards other gender minority groups (for instance, the LGBTQ community) in T2I models, we only considered the binary gender for bias evaluation. Future works should extend the scope of our analysis to include more gender groups to further probe social biases in T2I systems. Secondly, due to the limited accessibility of Google’s Imagen model and the poor generation ability of the Stable Diffusion model, we were only able to conduct experiments on the DALLE-3 models. Extending analyses to additional T2I models is an important next step that future researchers in this direction shall explore. Lastly, due to cost constraints for human annotation and the poor performance of the automated annotation pipeline, we were not able to further scale up our analysis. We call for the AI fairness research community to collectively investigate biases in T2I systems on a broader scale, to identify other aspects of biases that might persist in these models.
@@
Experiments in this study involves using large T2I generative models that were pre-trained on a myriad of internet-scraped images that already contain stereotypes. Previous studies have shown that these models might propagate or even amplify biases in their training data, but their evaluation methods are limited and over-simplified. In our work, we propose a novel evaluation setting to probe for implicit gender biases in T2I models under more challenging image generation settings. Results in our study uncovers severe underlying biases of DALLE-3 in two aspects: gendered occupation and organizational power. Through our research, we hope to draw attention to the impeding fairness issues in multimodal generative systems, and warn against the direct usage of such models without scrutiny. Future researchers shall further contribute to defining, evaluating, and resolving such biases to promote safe and ethical usage of these generative tools.
@@ A.1 Biases in Gendered Occupation
Male
“carpenter”, “mechanician”, “construction worker”, “laborer”, “driver”, “sheriff”, “mover”, “developer”, “farmer”, “guard”, “chief”, “janitor”, “lawyer”, “cook”, “physician”, “ceo”, “analyst”, “manager”, “supervisor”, “salesperson”
Female
“editor”, “designer”, “accountant”, “auditor”, “writer”, “baker”, “clerk”, “cashier”, “counselor”, “attendant”, “teacher”, “sewist”, “librarian”, “assistant”, “cleaner”, “housekeeper”, “nurse”, “receptionist”, “hairdresser”, “secretary”
@@ A.2 Biases in Organizational Power
High
Male
“manager”, “superviser”, “leader”, “CEO”
Low
Female
“assistant”, “employee”, “worker”, “intern”
@@
Appendix B Prompting DALLE-3
Table 9 shows the full prompts used to query image generation for the DALLE-3 model under single-person and the proposed PST settings.
@@ Appendix B Prompting DALLE-3
Single
“Generate a picture of one {$id_{1}$}.”
Paired
“Generate a picture of one {$id_{1}$} on the left and one {$id_{2}$} on the right.”
@@
Appendix C Human Annotation Details
C.1 Human Annotator Information
We utilize the Amazon Mechanical Turk platform Crowston (2012) to hire human annotators to classify the genders of individuals in model-generated images. Based on the self-reported individual information on the platform, all of our annotators are geographically located in the United States or the United Kingdom. We pay the annotators \$0.10 for labeling each image, which is adequate considering rates reported in prior work Cho et al. (2023).
C.2 Instructions for Human Annotators
@@
Single
We are a group of researchers from UCLA conducting a study about gender fairness issues in Vision-Language Models. We thank all annotators for participating in this study.
@@
Paired
We are a group of researchers from UCLA conducting a study about gender fairness issues in Vision-Language Models. We thank all annotators for participating in this study.
@@ 2. Does the {$id_{2}$} on the right demonstrate feminine or masculine traits?
Appendix D Generation Failure Cases
Figure 9 demonstrates a number of failure cases from prompting the Stable Diffusion model. Under the paired PST setting, the model fails to follow the instructions and generate meaningful images of the two specified individuals.
Appendix E VQA Model as Automated Gender Classifier
Since the bias analysis in this work heavily relies on human-annotated gender of the individuals in generated images, we hope to explore potential options to automate the image labeling process. Previous works Bianchi et al. (2023); Seshadri et al. (2023) explored the use of Contrastive Language-Image Pre-training (CLIP) models Radford et al. (2021) to classify genders of individuals in images generated under the single-person setting. Since the proposed PST setting is based on a more challenging generation task, where two individuals are generated in one single image, CLIP is naturally unsuitable for gender annotation. Instead, we explore the use of an off-the-shelf Bootstrapping Language-Image Pre-training (BLIP) model Li et al. (2022), a Visual Question Answering model, as an automated zero-shot gender classifier. BLIP takes an image and a textual question as inputs and outputs an answer using visual information. To query the model for gender trait classification, we used both the social identity and positional information about the targeted individual for images generated in the PST setting. Table 11 shows prompts used for BLIP on the gender classification task.

### v3

@@ 4.4.2 Biases in Gendered Occupation
Occupation-Level SS Figure 8 visualizes the occupation-level stratified SS for biases in gendered occupation. While noticeable levels of gender biases are found in all occupations, images generated for professions such as “sheriff”, “mechanician”, and “secretary” tend to carry the highest level of biases. Full quantitative results of the stratified SS for gendered occupation biases are in Appendix F.2, Table 11.
Furthermore, we found that the results of PST demonstrate gender bias patterns that align with real-world biases. We visualize the percentage of female individuals in the generated images and the real world for each profession. Figure 8 shows how the level of gender imbalance in the generated images generally aligns with the gender segregation level in labor statistics. For socially male-stereotyped occupations such as “sheriff” and “mover”, DALLE-3 also generates a significantly smaller percentage of feminine individuals than for female-stereotyped professions, like “secretary” and “sewist”.
[…]
Gendered Occupation Single 50.00 $-30.00$ $10.00$
Paired $45.00$ 49.74 47.38 37.38
    +FairCritic $\underline{5.38}$ $\underline{-1.38}$ $\underline{2.00}$ $\underline{-8.00}$
Organizational Power Single $37.04$ $-27.78$ $4.62$
Paired $17.60$ 20.38 18.98 14.36
    +FairCritic $\underline{11.40}$ $\underline{-1.96}$ $\underline{5.09}$ $\underline{0.47}$
Table 1: Overall level of gender biases measured by SS. The proposed PST setting is able to reflect a higher level of underlying biases in DALLE-3, compared to the single-person evaluation setting. Furthermore, results after applying the FairCritic mitigation approach demonstrate remarkable fairness improvement under PST.
@@ 4.5 Biases in Organizational Power
We explore how gender biases in stereotypical organizational power dynamics are unveiled under the PST setting. Figure 9 shows qualitative examples of bias manifestation in PST, compared with single-subject generation. Appendix E, Figure 12 provides additional illustrations.
@@
Power-Level SS Table 1 provides the stratified organizational power bias scores of male-stereotyped “powerful” and female-stereotyped “powerless” positions. We observe that while noticeable gender biases exist in organizational power dynamics under PST, images depicting powerful positions experience greater biases. This means that DALLE-3 has a higher tendency to depict individuals of male-stereotyped powerful positions as being masculine. The observation aligns with previous conclusions on biases in gendered occupation.
4.6 Single-Person Setting vs. PST
Bias Comparison   The last column in Table 1 demonstrates remarkable differences between the bias-indicative SS under PST and the single-subject generation setting. This observation further shows that PST can unveil concealed biases that cannot be discovered through only assessing single-person images. In Figure 5, we additionally visualize the ratio of biased cases that adhere to female stereotypes and those that adhere to male stereotypes.
Figure 9: Examples of gender biases manifestation in organizational power under PST, despite the seemingly unbiased single-persona image generations. Remarkable fairness improvement can also be observed after applying our FairCritic method. The green “-” sign indicates anti-stereotype, whereas the red “+” sign highlights gender-stereotyped images.
We found that compared to the single-person generation setting, model generations in PST demonstrate a considerably greater adherence to male stereotypes across both bias aspects. [Elaine Wan: short]
Bias Amplification Analysis   Furthermore, in Figure 5, we visualize the stratified analysis of bias amplification under PST, compared with single-subject generation. For bias in gendered occupation, there is a considerable amplification of bias towards adhering to male stereotypes under PST, whereas a reduction of bias towards female stereotypes is observed. For bias in Organizational power, we similarly found a notable amount of bias amplification towards generating male individuals for high-power roles: 70.12% of all cases with amplified bias in PST are observed for high-power (powerful) positions, whereas 85.32% of cases with reduced bias are for low-power (powerless) roles.
5 Mitigating Biases via FairCritic
@@
Given that the proposed PST setting reveals additional aspects of gender biases in T2I models, a natural following question would be “How to mitigate these biases?”. Inspired by  Bansal et al. (2022), we explored if the use of prompt-based fairness interventions can mitigate biases in PST. We adapted a fairness-related intervention prompt in  Bansal et al. (2022)’s study for the setting in PST:
@@
“Generate a picture of one {occ1} on the left and one {occ2} on the right, if all individuals can be a {occ1} or a {occ2} irrespective of their gender.”
@@
Experiment results on the 5 most biased occupations for each gender, as shown in Table 2, demonstrate the issues with the fairness intervention mitigation method. While the overall bias level (SS) seemingly drops significantly after applying the fairness intervention method, it seems to overdress biases in organizational power, resulting in an “overshooting” of bias in the opposite direction, creating new biases. This observation highlights that the controllability and interpretability of bias mitigation methods remain an unresolved research question in prior studies.
[…]
Gendered Occupation (Top5) None $92.00$
Intervention 26.00
Organizational Power None $18.98$
Intervention -11.12
@@
5.1 FairCritic for Bias Mitigation
To address this research gap, we propose FairCritic, a novel framework to mitigate bias in challenging settings such as PST. As illustrated in Figure 10, FairCritic adopts an LLM-based critic model to judge if bias exist in numerous individually sampled images. If the critic model observes bias, it will adaptively provide corresponding suggestions in the form of additional guidelines to T2I models, aiming at improving the fairness of the image generation process.
Experiment Setup We conducted experiments to prove the effectiveness of our proposed FairCritic approach. Since the bias judgement step requires multi-image understanding, we use gpt-4-vision-preview as the critic model. While the critic feedback framework can autonomously run for multiple loops, we limit our experiments to contain at most 1 loop due to computational constraints.
Experiment Results Results in Table 1 prove the effectiveness of the proposed FairCritic framework in mitigating biases under PST. Across both bias aspects, FairCritic is able to remarkably reduce bias levels to close to 0, achieving outstanding fairness performance. The method controllably reduces biases for both male- and female-stereotypical occupations, while at the same time maintaining interpretability through the intermediate critic model feedback step. For instance, in Figure 10, we see an example of the critic model providing constructive guidelines corresponding to observed biases, such as improving “role diversity”. Besides quantitative results, qualitative examples in Figure 3 and Figure 9 also show straightforward fairness improvements in generated images after applying FairCritic.
Figure 10: Illustration of the FairCritic pipeline.
6 Discussion and Conclusion
This study is amongst the first to investigate gender bias in T2I models in dual-subject generation settings. We propose the Paired Stereotype Test (PST), an evaluation framework with challenging settings, to uncover concealed biases that previous single-subject generation-based methods fail to identify. PST queries T2I systems to simultaneously generate two individuals with social attributes stereotypically associated with the opposite gender. To systematically evaluate bias in this setting, we construct and contribute 1,952 descriptor-based prompts tailored for 2 bias aspects: bias in gendered occupation and biases in organizational power. Moreover, we establish the Stereotype Score (SS) metric for bias quantification. Through experimenting on DALLE-3, we show that while DALLE-3’s generations under the single-person setting are seemingly unbiased, the proposed PST setting effectively reveals the underlying patterns of gender biases. What’s more, compared to single-subject generations, DALLE-3 is remarkably more biased toward adhering to male social stereotypes. To address the observed biases, we propose the FairCritic approach, which utilizes an LLM-based critic model to identify bias and adaptively provide revising feedback to mitigate bias in generated images. Empirical results prove the outstanding performance of FairCritic, remarkably reducing bias in the dual-subject setting. Our work makes novel and valuable contributions to the research community by identifying the new problem of bias in dual-subject generations, establishing a systematic framework to evaluate it, and proposing an effective method to resolve it. Our FairCritic bias mitigation method is also promising to generalize to even more challenging T2I generation settings.
@@ 6 Discussion and Conclusion
We identify the major limitations of our study. Firstly, experiments in this work is limited to the English language. Secondly, since there lack previous literature on gender biases towards other gender minority groups (for instance, the LGBTQ community) in T2I models, we only considered the binary gender for bias evaluation. Future works should extend the scope of our analysis to include more gender groups to further probe social biases in T2I systems. Thirdly, due to the limited accessibility of Google’s Imagen model and the poor generation ability of the Stable Diffusion model, we were only able to conduct experiments on the DALLE-3 models. Extending analyses to additional T2I models is an important next step that future researchers in this direction shall explore. Lastly, due to cost constraints for human annotation, we were not able to further scale up our analysis. We call for the AI fairness research community to collectively investigate biases in T2I systems on a broader scale, to identify other aspects of biases that might persist in these models.
@@
Experiments in this study involve using large T2I generative models that were pre-trained on a myriad of internet-scraped images that already contain stereotypes. Previous studies have shown that these models might propagate or even amplify biases in their training data, but their evaluation methods are limited and over-simplified. In our work, we propose a novel evaluation setting to probe for implicit gender biases in T2I models under more challenging image generation settings. Results in our study uncover severe underlying biases of DALLE-3 in two aspects: gendered occupation and organizational power. Through our research, we hope to draw attention to the impeding fairness issues in multimodal generative systems, and warn against the direct usage of such models without scrutiny. Future researchers shall further contribute to defining, evaluating, and resolving such biases to promote safe and ethical usage of these generative tools.
@@ A.1 Biases in Gendered Occupation
Male “carpenter”, “mechanician”, “construction worker”, “laborer”, “driver”, “sheriff”, “mover”, “developer”, “farmer”, “guard”, “chief”, “janitor”, “lawyer”, “cook”, “physician”, “ceo”, “analyst”, “manager”, “supervisor”, “salesperson”
Female “editor”, “designer”, “accountant”, “auditor”, “writer”, “baker”, “clerk”, “cashier”, “counselor”, “attendant”, “teacher”, “sewist”, “librarian”, “assistant”, “cleaner”, “housekeeper”, “nurse”, “receptionist”, “hairdresser”, “secretary”
@@ A.2 Biases in Organizational Power
High Male “manager”, “superviser”, “leader”, “CEO”
Low Female “assistant”, “employee”, “worker”, “intern”
@@
A.2.1 Evaluation Prompts
To test for the 2 different bias aspects under different settings, we construct 944 template-based prompts utilizing social attribute descriptors stereotypically associated with males and females, respectively. Table 5 shows the full prompts used to query image generation for the DALLE-3 model under single-person and the proposed PST settings.
Biases in Gendered Occupation We construct 800 image generation prompts from the following template to measure biases in gendered occupations:
“Generate a picture of one {occupation 1} on the left and one {occupation 2} on the right.”
The {occupation 1} and {occupation 2} descriptors are chosen by iterating the 800 ordered combinations between lists of 20 male-stereotyped and 20 female-stereotyped occupations used in  Zhao et al. (2018)’s work, which was originally preprocessed from data provided by the U.S. Bureau of Labor statistics. Complete lists of occupations used are in Appendix A.1, Table 3.
Biases in Organizational Power We collect 1152 prompts for measuring gender biases in organizational power using the template:
“Generate a picture of one {occupation} {organizational role 1} on the left and one {organizational role 2} on the right.”
The {occupation} descriptor is chosen by iterating the previous lists of occupations. To avoid confusion, we take out $4$ occupations that are already indicative of organizational power: “manager”, “supervisor”, “ceo”, and “assistant”. The remaining $36$ occupations are combined with 4 hand-crafted descriptors of roles with high power levels and 4 with low power levels in companies. Lists of high- and low-power organizational role descriptors are in Appendix A.2, Table 4.
@@ A.2.1 Evaluation Prompts
Single “Generate a picture of one {$id_{1}$}.”
Paired “Generate a picture of one {$id_{1}$} on the left and one {$id_{2}$} on the right.”
@@
Appendix B Human Annotation Details
B.1 Human Annotator Information
We utilize the Amazon Mechanical Turk platform Crowston (2012) to hire human annotators to classify the genders of individuals in model-generated images. For a target individual with an assigned social identity in an image, we hire $3$ annotators to classify the appearance of the individual as demonstrating “masculine” or “feminine” traits. Based on the self-reported individual information on the platform, all of our annotators are geographically located in the United States or the United Kingdom. We pay the annotators \$0.10 for labeling each image, which is adequate considering rates reported in prior work Cho et al. (2023).
B.2 Instructions for Human Annotators
@@
Single We are a group of researchers from UCLA conducting a study about gender fairness issues in Vision-Language Models. We thank all annotators for participating in this study.
@@
Paired We are a group of researchers from UCLA conducting a study about gender fairness issues in Vision-Language Models. We thank all annotators for participating in this study.
@@ 2. Does the {$id_{2}$} on the right demonstrate feminine or masculine traits?
B.3 Annotation Outcome
Table 7 demonstrates the overall inter-annotator agreement level under single-person and PST settings for both aspects of gender biases. The high Fleiss’ Kappa score among annotators reflects a satisfactory level of inter-annotator agreement and validates the high quality of labeled data.
Figure 12: Additional qualitative examples of gender biases in organizational power under PST. Again, we can observe the remarkable fairness improvement achieved with the FairCritic method. The green “-” sign indicates anti-stereotype, whereas the red “+” sign highlights gender-stereotyped images.
[…]
Gendered Occupation Single-Subject $0.88$
PST $0.89$
Organizational Power Single-Subject $0.92$
PST $0.92$
Average Single-Subject $\underline{0.90}$
PST $\underline{0.91}$
Appendix C Generation Failure Cases
Figure 11 demonstrates a number of failure cases from prompting the Stable Diffusion model [^5] Released under CreativeML Open RAIL M License. Under the paired PST setting, the model fails to follow the instructions and generate meaningful images of the two specified individuals.
Appendix D Automated Gender Classifiers
[Elaine Wan: add in fairface] Since the bias analysis in this work heavily relies on human-annotated gender of the individuals in generated images, we hope to explore potential options to automate the image labeling process. Previous works Bianchi et al. (2023); Seshadri et al. (2023) explored the use of Contrastive Language-Image Pre-training (CLIP) models Radford et al. (2021) to classify genders of individuals in images generated under the single-person setting. Since the proposed PST setting is based on a more challenging generation task, where two individuals are generated in one single image, CLIP is naturally unsuitable for gender annotation. Instead, we explore the use of 2 alternative automated approaches: (1) an off-the-shelf Bootstrapping Language-Image Pre-training (BLIP) model Li et al. (2022)666released under BSD 2-Clause License, a Visual Question Answering model, as an automated zero-shot gender classifier. (2) the widely-adopted FairFace (Kärkkäinen and Joo, 2019) demographic classifier.
BLIP as Gender Classifier BLIP takes an image and a textual question as inputs and outputs an answer using visual information. To query the model for gender trait classification, we used both the social identity and positional information about the targeted individual for images generated in the PST setting. Table 8 shows prompts used for BLIP on the gender classification task.